How long does it reach for a particle to reach a distance of 100 km If its average velocity is 100 m/s


1
Expert's answer
2021-11-17T10:38:09-0500
"t=\dfrac{d}{v}=\dfrac{10^5\ m}{100\ \dfrac{m}{s}}=1000\ s\times\dfrac{1\ h}{3600\ s}=0.28\ h."